OpenStudy (anonymous):

Which point lies outside the circle with equation (x - 2)^2 + (y + 3)^2 = 4?

OpenStudy (anonymous):

@Ahmad1

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

where the choices are dude @cocopuffs90 ?

OpenStudy (anonymous):

a) 2,4 b) 2,-4 c) 0,-3 d) 1,-4

OpenStudy (anonymous):

@Ahmad1

OpenStudy (anonymous):

okay in these types of questions , all you have to do is just substititute the points instead of (x,y) if the equation is not satisfied then that's the answer.. try and tell me what you see!

OpenStudy (anonymous):

is it b?

OpenStudy (anonymous):

@Ahmad1

OpenStudy (anonymous):

@ybarrap do you happen to know if its b?

OpenStudy (ybarrap):

Compute the distance for each or your points to the center of the circle (<2,-3>). If it's greater than 2, that's your answer.

OpenStudy (anonymous):

im still confused... so its not b right? @ybarrap

OpenStudy (anonymous):

I think you are right dude @cocopuffs90

OpenStudy (ybarrap):

I just computed the distances, took me less than a minute and the answer just pops out. Do you know the distance between points? It's not b.

OpenStudy (anonymous):

thanks man @Ahmad1

OpenStudy (anonymous):

oh im wrong... and lost. i dont know the distance @ybarrap

OpenStudy (anonymous):

oooooooooooooooops I am so sorry @cocopuffs90

OpenStudy (anonymous):

@Ahmad1 Its all good dude

OpenStudy (anonymous):

I did not read the question properly , what a shame!

OpenStudy (anonymous):

haha its okay. so im guessing its (2,0) then right?

OpenStudy (ybarrap):

$$\sqrt{(y_2 - y_1)^2+(x_2 - x_1)^2}$$

OpenStudy (anonymous):

even though it's right dude @cocopuffs90 the idea of the question is that the distance between the center and the point must be bigger than the radius , I also think tha's what our friend @ybarrap is saying

OpenStudy (anonymous):

the center is the point (2,-3)

OpenStudy (ybarrap):

For example, for your first point, the distance is \(|<2,-3> - <2,4>|=\sqrt{0^2+(-7)^2}=7\) All the others are just as easy.

OpenStudy (anonymous):

but (2,-3) isnt an option..

OpenStudy (ybarrap):

That's the center of the circle. This formula determines the distance from this center to each point in your options.

OpenStudy (anonymous):

oooooohh so what exactly do i do now

OpenStudy (anonymous):

I know but the distgance is measured between (2,-3) and the other points that distance must be bigger than the radius which is 2

OpenStudy (anonymous):

im not sure i know how to do that.. the formula

OpenStudy (ybarrap):

You need to find the distance for the next point, from <2,-4> to the center.

OpenStudy (anonymous):



OpenStudy (ybarrap):